OpenStudy (anonymous):

A cooler at a picnic contains 100 sodas covered by ice. There are 11 cans of cola, 28 cans of orange juice, 15 of ginger rale, and 46 of rootbear. these cans are all in the same shapes and sizes. if one can is selected at ramdon what is the probl that the soda is ginger ale, rootbear or cola?

OpenStudy (amistre64):

add the three totals and slap a % to it

OpenStudy (amistre64):

the probablitity of a gingerbearcola is....

OpenStudy (anonymous):

89?

OpenStudy (anonymous):

all toghether

OpenStudy (amistre64):

72 is what i count: 11,46,15

OpenStudy (anonymous):

so u just add them toghether for the answer

OpenStudy (amistre64):

the converse is easier to determine 100 - 28 = 72

OpenStudy (amistre64):

yes; you wanna know what is the chance that youl get any of those 3

OpenStudy (amistre64):

since there was 100 to choose from; it makes it easier to just add the number

OpenStudy (anonymous):

72 becomes the bottom number

OpenStudy (anonymous):

now do u subtract 72 from 100

OpenStudy (amistre64):

72 out of 100 72/100

OpenStudy (anonymous):

and to reduce that you device them

OpenStudy (amistre64):

28 cans are not gingerbearcola sooo 1 - 28/100 is the same answer as 72/100

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

36/50 = 18/25 ...

OpenStudy (anonymous):

and keep reducin

OpenStudy (anonymous):

or 18/25 is lowest

OpenStudy (anonymous):

3/5?